Women have pulse rates that are normally distributed with a mean of
80.280.2
beats per minute and a standard deviation of
10.910.9
beats per minute. Complete parts a through c below.

a. Find the percentiles
Upper P 1P1
and
Upper P 99P99.

Answer by Boreal(15235)

The 1% percentile is z=-2.328, the 99th z=+2.328
z*sd=2.328*10.910=25.398
lower is 54.882 bpm
upper is 105.678 bpm